# Nearest-Neighbours Neural Network architecture for efficient sampling of statistical physics models

### v1

@@ Abstract
The task of sampling efficiently the Gibbs-Boltzmann distribution of disordered systems is important both for the theoretical understanding of these models and for the solution of practical optimization problems. Unfortunately, this task is known to be hard, especially for spin glasses at low temperatures. Recently, many attempts have been made to tackle the problem by mixing classical Monte Carlo schemes with newly devised Neural Networks that learn to propose smart moves. In this article we introduce the Nearest-Neighbours Neural Network (4N) architecture, a physically-interpretable deep architecture whose number of parameters scales linearly with the size of the system and that can be applied to a large variety of topologies. We show that the 4N architecture can accurately learn the Gibbs-Boltzmann distribution for the two-dimensional Edwards-Anderson model, and specifically for some of its most difficult instances. In particular, it captures properties such as the energy, the correlation function and the overlap probability distribution. Finally, we show that the 4N performance increases with the number of layers, in a way that clearly connects to the correlation length of the system, thus providing a simple and interpretable criterion to choose the optimal depth.
@@ Introduction
parametrized by ‘local fields’ $H_{i}$ and ‘pairwise couplings’ $J_{ij}$. In physics applications such as spin glasses, these are often chosen to be i.i.d. random variables, e.g. $H_{i}\sim\mathcal{N}(0,H^{2})$ and $J_{ij}\sim\mathcal{N}(0,J^{2}/N)$. In Boltzmann Machines, instead, the fields and couplings are learned by maximizing the likelihood of a given training set.
In many cases, dealing with Eq. (1) analytically, e.g. computing expectation values of quantities of interest, is impossible, and one resorts to numerical computations. A universal strategy is to use local Markov Chain Monte Carlo (MCMC) methods, which have the advantage of being applicable to a wide range of different systems. In these methods, one proposes a local move, typically flipping a single spin, $\sigma_{i}\to-\sigma_{i}$, and accepts or rejects the move in such a way to guarantee that after many iterations, the configuration $\boldsymbol{\sigma}$ is distributed according to Eq. (1). Unfortunately, these methods are difficult (if not altogether impossible) to apply in many hard-to-sample problems and, in practice, they may end up taking a huge amount of time. For instance, finding the ground state of the Sherrington-Kirkpatrick model (a particular case of sampling at $T\to 0$) is known to be a NP-hard problem; sampling a class of optimization problems is known to take a time scaling exponentially with $N$. The training of Boltzmann Machines is also known to suffer from this kind of problems.
@@
Because $\mathcal{H}$ can be computed in polynomial time and $P_{\text{NN}}$ can be sampled efficiently by hypothesis, this approach ensures an efficient sampling of the Gibbs-Boltzmann probability, as long as (i) $P_{\text{NN}}$ covers well the support of $P_{\text{GB}}$ (i.e., the so-called mode collapse is avoided) and (ii) the acceptance probability in Eq. (5) is significantly different from zero for most of the generated configurations.
@@
The main problem is how to train the Neural Network to minimize Eq. (4); in fact, the computation of $D_{\rm KL}$ requires sampling from $P_{\text{GB}}$. A possible solution is to use $D_{\text{KL}}(P_{\text{NN}}\parallel P_{\text{GB}})$ instead [10], but this has been shown to be prone to mode collapse. Another proposed solution is to set up an iterative procedure, called ‘Sequential Tempering’ (ST), in which one first learns $P_{\text{NN}}$ at high temperature where sampling from $P_{\text{GB}}$ is possible, then gradually decreases the temperature, updating $P_{\text{NN}}$ slowly [11]. It is still unclear whether this kind of strategies actually perform well on hard problems and challenge already-existing algorithms, with both positive [12, 13, 14] and negative results [15, 16, 17, 18]. First steps in a theoretical understanding of the performances of such algorithms are just being made [19]. A variety of new methods and techniques have thus been proposed to tackle the problem, such as autoregressive models [10, 11], normalizing flows [20, 21, 22, 23], and diffusion models [24]. The main drawbacks of these architectures is that the number of parameters scales poorly with the system size, typically as $N^{\alpha}$ with $\alpha>1$, while local Monte Carlo requires a number of operations scaling linearly in $N$. Furthermore, these architectures – with the notable exception of TwoBo [25] that directly inspired our work – are not informed about the physical properties of the model, and in particular about the structure of its correlations.
@@
In this paper, we introduce the Nearest Neighbours Neural Network, or 4N for short, a Graph Neural Network-inspired architecture than implements an autoregressive scheme to learn the Gibbs-Boltzmann distribution. This architecture has a number of parameters linear in the system size and can sample new configurations in $\mathcal{O}(N)$ time, thus achieving the best possible scaling one could hope for such architectures. Moreover, 4N has a straightforward physical interpretation. In particular, the choice of the number of layers can be directly linked to the correlation length of the model under study. Finally, the architecture can easily be applied to essentially any statistical system, such as lattices in higher dimensions or random graphs, and it is thus more general than other architectures. As a proof of concept, we evaluate the 4N architecture on the two-dimensional Edwards-Anderson spin glass model. We demonstrate that the model succeeds in accurately learning the Gibbs-Boltzmann distribution, especially for some of the most challenging model instances. Notably, it precisely captures properties such as energy, the correlation function, and the overlap probability distribution.
@@
where $\boldsymbol{\sigma}_{<i}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{i-1})$, so that states can then be generated using ancestral sampling, i.e. generating first $\sigma_{1}$, then $\sigma_{2}$ conditioned on the sampled value of $\sigma_{1}$, then $\sigma_{3}$ conditioned to $\sigma_{1}$ and $\sigma_{2}$ and so on. The factorization in (6) is exact, but computing $P(\sigma_{i}\mid\boldsymbol{\sigma}_{<i})$ exactly generally requires a number of operations scaling exponentially with $N$. Hence, in practice, $P_{\text{GB}}$ is approximated by $P_{\text{NN}}$ that takes the form in Eq. (6) where each individual term $P(\sigma_{i}\mid\boldsymbol{\sigma}_{<i})$ is approximated using a small set of parameters. Note that the unsupervised problem of approximating $P_{\text{GB}}$ is now formally translated into a supervised problem of learning the output, i.e. the probability $\pi_{i}\equiv P(\sigma_{i}=+1\mid\boldsymbol{\sigma}_{<i})$, as a function of the input $\boldsymbol{\sigma}_{<i}$. The specific way in which this approximation is carried out depends on the architecture. In this section we will describe some common autoregressive architectures found in literature, for approximating the function $\pi_{i}(\boldsymbol{\sigma}_{<i})$.
@@
    where $\mathbf{y}_{i}\in\mathbb{R}^{N_{h}}$, $\underline{\underline{A}}\in\mathbb{R}^{N_{h}\times N}$ and $\boldsymbol{\sigma}_{<i}$ is the vector of the spins in which the spins $l\geq i$ have been masked, i.e. $\mathbf{\sigma}_{<i}=(\sigma_{1},\sigma_{2},\dots,\sigma_{i-1},0,\dots,0)$. $\mathbf{B}\in\mathbb{R}^{N_{h}}$ is the bias vector and $\Psi$ is an element-wise activation function. Note that $A$ and $\mathbf{B}$ do not depend on the index $i$, i.e. they are shared across all spins. The information from the hidden layer is then passed to a fully connected layer
    $H_{i}=\Psi\left(\mathbf{V}_{i}\cdot\mathbf{y}_{i}+C_{i}\right)$ (8)
    where $\mathbf{V}_{i}\in\mathbb{R}^{N_{h}}$ and $C_{i}\in\mathbb{R}$. Finally, the output probability is obtained by applying a sigmoid function to the output, $\pi_{i}=S(H_{i})$. The Nade Architecture uses $\mathcal{O}(N\times N_{h})$ parameters. However, the choice of the hidden dimension $N_{h}$ and how it should be related to $N$ is not obvious. For this architecture to work well in practical applications, $N_{h}$ needs to be $\mathcal{O}(N)$, thus leading to $\mathcal{O}(N^{2})$ parameters [11].
@@
  • Autoregressive Convolutional Neural Network (CNN) architectures, such as PixelCNN [27], are networks that implement a CNN structure but superimpose the autoregressive property. These architecture typically use translationally invariant kernels ill-suited to study disordered models and remain limited to two-dimensional systems.
@@
    Therefore, the TwoBo architecture first computes the vector $\boldsymbol{h}^{i}$, then approximates $\rho_{i}(\boldsymbol{h}^{i})$ using a Neural Network (for instance, a Multilayer Perceptron) and finally computes Eq. (9) using a skip connection (for the term $2\beta h^{i}_{i}$) and a sigmoid activation. By construction, in a two-dimensional model with $N=L^{2}$ variables and nearest-neighbor interactions, for given $i$ only $\mathcal{O}(L=\sqrt{N})$ of the $h^{i}_{l}$ are non-zero, see Ref. [25, Fig.1]. Therefore, the TwoBo architecture has $\mathcal{O}(N^{\frac{3}{2}})$ parameters in the two-dimensional case. However, the scaling becomes worse in higher dimensions $d$, i.e. $\mathcal{O}(N^{\frac{2d-1}{d}})$ and becomes the same as MADE (i.e. $\mathcal{O}(N^{2})$) for random graphs. Additionally, the choice of the $h^{i}_{l}$, although justified mathematically, does not have a clear physical interpretation: the TwoBo architecture takes into account far away spins which, however, should matter much when $N$ is large.
To summarize, the NADE, MADE and TwoBo architectures all need $\mathcal{O}(N^{2})$ parameters for generic models defined on random graph structures. TwoBo has less parameters for models defined on $d$-dimensional lattices, but still $\mathcal{O}(N^{\frac{2d-1}{d}})\gg\mathcal{O}(N)$. CNN architectures with translationally invariant kernels have less parameters, but are not suitable for disordered systems and remain limited to $d=2$.
@@ Results
The 4N architecture (Fig. 1) computes the probability $\pi_{i}\equiv P(\sigma_{i}=+1\mid\boldsymbol{\sigma}_{<i})$ by propagating the $h_{l}^{i}$, defined as in TwoBo, Eq. (10), through a Graph Neural Network (GNN) architecture. The variable $h_{l}^{i}$ is interpreted as the local field induced by the frozen variables $\boldsymbol{\sigma}_{<i}$ on spin $l$. Note that in 4N, we consider all possible values of $l=1,\cdots,N$, not only $l\geq i$ as in TwoBo. The crucial difference with TwoBo is that only the fields in a finite neighborhood of variable $i$ need to be considered, because the initial values of the fields are propagated through a series of GNN layers that take into account the locality of the physical problem. During propagation, fields are weighted using layer and edge dependent layers together with the couplings $J_{ij}$ and the inverse temperature $\beta$. After each layer of the network (except the final one), a skip connection to the initial configuration is added. After the final layer, a sigmoidal activation is applied to find $\pi_{i}$. Details are given in the Methods section.
More precisely, because GNN updates are local, and there are $\ell$ of them, in order to compute the final field on $i$ we only need to consider fields at distance at most $\ell$. Therefore, the number of computations required to compute $\pi_{i}$ is $\mathcal{O}(\ell^{d})$, thus granting the generation of a new configuration in $\mathcal{O}(\ell^{d}\times N)$ steps. It is therefore crucial to asses how the number of layers $\ell$ should scale with the system size $N$.
@@
In this paper, in order to test the 4N architecture on a prototypically hard-to-sample problem, we will consider the Edwards-Anderson spin glass model, described by the Hamiltonian in Eq. (3) for Ising spins, where we set $H_{i}=0$ for simplicity. The couplings are chosen to be non-zero only for neighboring pairs $\langle i,j\rangle$ on a two-dimensional hypercubic lattice. Non-zero couplings $J_{ij}$ are independent random variables, identically distributed according to either a normal or a Rademacher distribution. This model was also used as a benchmark in Refs. [11, 15].
While this model has no finite-temperature spin glass transition, sampling it via standard local MCMC becomes extremely hard for $\beta\gtrsim 1.5$ [15]. To this day, the best option for sampling this model at lower temperature is parallel tempering (PT), which we used to construct a sample of $M=2^{16}$ equilibrium configurations for several instances of the model and at several temperatures. Details are given in the Methods section.
@@
We begin by comparing the mean energy of the configurations generated by the model with the mean energy of the configurations generated via parallel tempering, which is important because the energy plays an important role in the Metropolis reweighing scheme of Eq. (5). Fig. 2 shows the relative energy difference, as a function of inverse temperature $\beta$ and number of layers $\ell$, averaged over 10 different instances of the disorder. First, it is clear that adding layers improves results remarkably. We stress that, since the activation functions of our network are linear, adding layers only helps in taking into account spins that are further away. Second, we notice that the relative error found when using 6 layers or more is small for all temperatures considered. Finally, we notice that the behaviour of the error is non-monotonic, and indeed appears to decrease for low temperatures even for networks with few layers.
@@
Figure 4: Examples of the probability distribution of the overlap, $P(q)$, for different instances of the disorder and different temperatures. Solid color lines correspond to configurations generated using 4N, while dashed black lines correspond to the $P(q)$ obtained from equilibrium configurations generated via PT. (a) At high temperature, even a couple of layers are enough to reproduce well the distribution. (b) At lower temperatures more layers are required, as expected because of the increase of the correlation length and of the complexity of the Gibbs-Boltzmann distribution. (c) Even lower temperature for the same instance. (d) A test on a particularly complex instance shows that 4N is capable to learn even highly non-trivial forms of $P(q)$.
@@
Figure 6: Comparison between the correlation length $\xi$ and the decay parameter $\overline{\ell}$ of the exponential fit as a function of the inverse temperature $\beta$. Both data sets are averaged over 10 instances. The data are limited to $T\geq 0.58$ because below that temperature the finite size of the sample prevents us to obtain a reliable proper estimate of the correlation length.
@@
$G(r)=\frac{1}{L^{2}\Omega(r)}\sum_{i,j:d_{ij}=r}\overline{\langle q_{i}q_{j}\rangle}\ ,$ (13)
where $r=d_{ij}$ is the lattice distance of sites $i,j$, ${q_{i}=\sigma^{1}_{i}\sigma^{2}_{i}}$ is the local overlap, and $\Omega(r)=\sum_{i:d_{ij}=r}1$ is the volume of a ball of radius $r$ around any given site. Note that in this way $G(0)=1$ by construction. Results for $G(r)$ at various temperatures are shown in Fig. 5. It is clear that increasing the number of layers leads to a better agreement at lower temperatures. Interestingly, architectures with a lower numbers of layers produce a smaller correlation function at any distance, which shows that upon increasing the number of layers the 4N architecture is effectively learning the correlation of the Gibbs-Boltzmann measure.
To quantify this intuition, we extract from the exact $G(r)$ of the EA model a correlation length $\xi$, from the relation $G(\xi)=1/e$. Then, we compare $\xi$ with $\overline{\ell}$, i.e. the number of layers above which the model becomes accurate, extracted from Fig. 3. These two quantities are compared in Fig. 6 as a function of temperature, showing that they are roughly proportional. We are thus able to introduce a physical criterion to fix the optimal number of layers needed for the 4N architecture to perform well: $\ell$ must be a multiple of the correlation length $\xi$ of the model at that temperature. We conclude that, for a model with a correlation length $\xi$ that remains finite for $N\to\infty$, the 4N architecture requires $(c+1)\times\ell\times N$ parameters and $\mathcal{O}(\ell^{d}\times N)$ operations to generate a new configuration, with $\ell$ proportional to $\xi$, hence achieving the desired $\mathcal{O}(N)$ scaling with system size both for the number of parameters and the number of operations.

### v2

@@ Abstract
The task of sampling efficiently the Gibbs-Boltzmann distribution of disordered systems is important both for the theoretical understanding of these models and for the solution of practical optimization problems. Unfortunately, this task is known to be hard, especially for spin-glass-like problems at low temperatures. Recently, many attempts have been made to tackle the problem by mixing classical Monte Carlo schemes with newly devised Neural Networks that learn to propose smart moves. In this article, we introduce the Nearest-Neighbours Neural Network (4N) architecture, a physically interpretable deep architecture whose number of parameters scales linearly with the size of the system and that can be applied to a large variety of topologies. We show that the 4N architecture can accurately learn the Gibbs-Boltzmann distribution for a prototypical spin-glass model, the two-dimensional Edwards-Anderson model, and specifically for some of its most difficult instances. In particular, it captures properties such as the energy, the correlation function and the overlap probability distribution. Finally, we show that the 4N performance increases with the number of layers, in a way that clearly connects to the correlation length of the system, thus providing a simple and interpretable criterion to choose the optimal depth.
@@ Introduction
parametrized by ‘local external fields’ $H_{i}$ and ‘pairwise couplings’ $J_{ij}$. In physics applications such as spin glasses, these are often chosen to be independent and identically distributed random variables, e.g. $H_{i}\sim\mathcal{N}(0,H^{2})$ and $J_{ij}\sim\mathcal{N}(0,J^{2}/N)$. In Boltzmann Machines, instead, the fields and couplings are learned by maximizing the likelihood of a given training set.
In many cases, dealing with Eq. (1) analytically, e.g. computing expectation values of interesting quantities, is impossible, and one resorts to numerical computations. A universal strategy is to use local Markov Chain Monte Carlo (MCMC) methods, which have the advantage of being applicable to a wide range of different systems. In these methods, one proposes a local move, typically flipping a single spin, $\sigma_{i}\to-\sigma_{i}$, and accepts or rejects the move in such a way to guarantee that after many iterations, the configuration $\boldsymbol{\sigma}$ is distributed according to Eq. (1). Unfortunately, these methods are difficult (if not altogether impossible) to apply in many hard-to-sample problems for which the convergence time is very large, which, in practice, ends up requiring a huge amount of computational time. For instance, finding the ground state of the Sherrington-Kirkpatrick model (a particular case of sampling at $T\to 0$) is known to be a NP-hard problem; sampling a class of optimization problems is known to take a time scaling exponentially with $N$. The training of Boltzmann Machines is also known to suffer from this kind of problems.
@@
Because $\mathcal{H}$ can be computed in polynomial time and $P_{\text{NN}}$ can be sampled efficiently by hypothesis, this approach ensures an efficient sampling of the Gibbs-Boltzmann probability, as long as (i) $P_{\text{NN}}$ covers well the support of $P_{\text{GB}}$ (i.e., the so-called mode collapse is avoided) and (ii) the acceptance probability in Eq. (5) is significantly different from zero for most of the generated configurations. This scheme can also be combined with standard local Monte Carlo moves to improve efficiency [10].
@@
The main problem is how to train the Neural Network to minimize Eq. (4); in fact, the computation of $D_{\rm KL}$ requires sampling from $P_{\text{GB}}$. A possible solution is to use $D_{\text{KL}}(P_{\text{NN}}\parallel P_{\text{GB}})$ instead [11], but this has been shown to be prone to mode collapse [12]. Another proposed solution is to set up an iterative procedure, called ‘Sequential Tempering’ (ST), in which one first learns $P_{\text{NN}}$ at high temperature where sampling from $P_{\text{GB}}$ is possible, then gradually decreases the temperature, updating $P_{\text{NN}}$ slowly [13]. A variety of new methods and techniques have thus been proposed to tackle the problem, such as autoregressive models [11, 13, 14], normalizing flows [15, 10, 16, 17], and diffusion models [18]. It is still unclear whether this kind of strategies actually performs well on hard problems and challenges already-existing algorithms [12]. The first steps in a theoretical understanding of the performances of such algorithms are just being made [19, 20]. The main drawback of these architectures is that the number of parameters scales poorly with the system size, typically as $N^{\alpha}$ with $\alpha>1$, while local Monte Carlo requires a number of operations scaling linearly in $N$. Furthermore, these architectures – with the notable exception of TwoBo [21] that directly inspired our work – are not informed about the physical properties of the model, and in particular about the structure of its interaction graph and correlations. It should be noted that the related problem of finding the ground state of the system (which coincides with zero-temperature sampling) has been tackled using similar ideas, again with positive [22, 23, 24, 25, 26] and negative results [27, 28, 29, 30].
@@
In this paper, we introduce the Nearest Neighbours Neural Network, or 4N for short, a Graph Neural Network-inspired architecture that implements an autoregressive scheme to learn the Gibbs-Boltzmann distribution. This architecture has a number of parameters scaling linearly in the system size and can sample new configurations in $\mathcal{O}(N)$ time, thus achieving the best possible scaling one could hope for such architectures. Moreover, 4N has a straightforward physical interpretation. In particular, the choice of the number of layers can be directly linked to the correlation length of the model under study. Finally, the architecture can easily be applied to essentially any statistical system, such as lattices in higher dimensions or random graphs, and it is thus more general than other architectures. As a proof of concept, we evaluate the 4N architecture on the two-dimensional Edwards-Anderson spin glass model, a standard benchmark also used in previous work [13, 12]. We demonstrate that the model succeeds in accurately learning the Gibbs-Boltzmann distribution, especially for some of the most challenging model instances. Notably, it precisely captures properties such as energy, the correlation function, and the overlap probability distribution.
@@
where $\boldsymbol{\sigma}_{<i}=(\sigma_{1},\sigma_{2},\ldots,\sigma_{i-1})$, so that states can then be generated using ancestral sampling, i.e. generating first $\sigma_{1}$, then $\sigma_{2}$ conditioned on the sampled value of $\sigma_{1}$, then $\sigma_{3}$ conditioned to $\sigma_{1}$ and $\sigma_{2}$ and so on. The factorization in (6) is exact, but computing $P(\sigma_{i}\mid\boldsymbol{\sigma}_{<i})$ exactly generally requires a number of operations scaling exponentially with $N$. Hence, in practice, $P_{\text{GB}}$ is approximated by $P_{\text{NN}}$ that takes the form in Eq. (6) where each individual term $P(\sigma_{i}\mid\boldsymbol{\sigma}_{<i})$ is approximated using a small set of parameters. Note that the unsupervised problem of approximating $P_{\text{GB}}$ is now formally translated into a supervised problem of learning the output, i.e. the probability $\pi_{i}\equiv P(\sigma_{i}=+1\mid\boldsymbol{\sigma}_{<i})$, as a function of the input $\boldsymbol{\sigma}_{<i}$. The specific way in which this approximation is carried out depends on the architecture. In this section, we will describe some common autoregressive architectures found in literature, for approximating the function $\pi_{i}(\boldsymbol{\sigma}_{<i})$.
@@
    where $\mathbf{y}_{i}\in\mathbb{R}^{N_{h}}$, $\underline{\underline{A}}\in\mathbb{R}^{N_{h}\times N}$ and $\boldsymbol{\sigma}_{<i}$ is the vector of the spins in which the spins $l\geq i$ have been masked, i.e. $\mathbf{\sigma}_{<i}=(\sigma_{1},\sigma_{2},\dots,\sigma_{i-1},0,\dots,0)$. $\mathbf{B}\in\mathbb{R}^{N_{h}}$ is the bias vector and $\Psi$ is an element-wise activation function. Note that $\underline{\underline{A}}$ and $\mathbf{B}$ do not depend on the index $i$, i.e. they are shared across all spins. The information from the hidden layer is then passed to a fully connected layer
    ${\color[rgb]{0,0,0}\psi_{i}}=\Psi\left(\mathbf{V}_{i}\cdot\mathbf{y}_{i}+C_{i}\right)$ (8)
    where $\mathbf{V}_{i}\in\mathbb{R}^{N_{h}}$ and $C_{i}\in\mathbb{R}$. Finally, the output probability is obtained by applying a sigmoid function to the output, $\pi_{i}=S({\color[rgb]{0,0,0}\psi_{i}})$. The Nade Architecture uses $\mathcal{O}(N\times N_{h})$ parameters. However, the choice of the hidden dimension $N_{h}$ and how it should be related to $N$ is not obvious. For this architecture to work well in practical applications, $N_{h}$ needs to be $\mathcal{O}(N)$, thus leading to $\mathcal{O}(N^{2})$ parameters [13].
@@
  • Autoregressive Convolutional Neural Network (CNN) architectures, such as PixelCNN [32, 33], are networks that implement a CNN structure but superimpose the autoregressive property. These architectures typically use translationally invariant kernels ill-suited to study disordered models.
@@
    Therefore, the TwoBo architecture first computes the vector $\boldsymbol{h}^{i}$, then approximates $\rho_{i}(\boldsymbol{h}^{i})$ using a Neural Network (for instance, a Multilayer Perceptron) and finally computes Eq. (9) using a skip connection (for the term $2\beta h^{i}_{i}$) and a sigmoid activation. By construction, in a two-dimensional model with $N=L^{2}$ variables and nearest-neighbor interactions, for given $i$ only $\mathcal{O}(L=\sqrt{N})$ of the $h^{i}_{l}$ are non-zero, see Ref. [21, Fig.1]. Therefore, the TwoBo architecture has $\mathcal{O}(N^{\frac{3}{2}})$ parameters in the two-dimensional case. However, the scaling becomes worse in higher dimensions $d$, i.e. $\mathcal{O}(N^{\frac{2d-1}{d}})$ and becomes the same as MADE (i.e. $\mathcal{O}(N^{2})$) for random graphs. Additionally, the choice of the $h^{i}_{l}$, although justified mathematically, does not have a clear physical interpretation: the TwoBo architecture takes into account far away spins which, however, should not matter much when $N$ is large.
To summarize, the NADE, MADE and TwoBo architectures all need $\mathcal{O}(N^{2})$ parameters for generic models defined on random graph structures. TwoBo has less parameters for models defined on $d$-dimensional lattices, but still $\mathcal{O}(N^{\frac{2d-1}{d}})\gg\mathcal{O}(N)$. CNN architectures with translationally invariant kernels have fewer parameters, but are not suitable for disordered systems, are usually applied to $d=2$ and remain limited to $d\leq 3$.
@@ Results
The 4N architecture (Fig. 1) computes the probability $\pi_{i}\equiv P(\sigma_{i}=+1\mid\boldsymbol{\sigma}_{<i})$ by propagating the $h_{l}^{i}$, defined as in TwoBo, Eq. (10), through a Graph Neural Network (GNN) architecture. The variable $h_{l}^{i}$ is interpreted as the local field induced by the frozen variables $\boldsymbol{\sigma}_{<i}$ on spin $l$. Note that in 4N, we consider all possible values of $l=1,\cdots,N$, not only $l\geq i$ as in TwoBo. The crucial difference with TwoBo is that only the fields in a finite neighborhood of variable $i$ need to be considered, because the initial values of the fields are propagated through a series of GNN layers that take into account the locality of the physical problem. During propagation, fields are combined with layer- and edge-dependent weights together with the couplings $J_{ij}$ and the inverse temperature $\beta$. After each layer of the network (except the final one), a skip connection to the initial configuration is added. After the final layer, a sigmoidal activation is applied to find $\pi_{i}$. The network has a total number of parameters equal to $(c+1)\times\ell\times N$, where $c$ is the average number of neighbors for each spin. Details are given in the Methods section.
More precisely, because GNN updates are local, and there are $\ell$ of them, in order to compute the final field on site $i$, hence $\pi_{i}$, we only need to consider the set of fields corresponding to sites at distance at most $\ell$, hence the number of operations scales proportionally to the number of such sites which we call $\mathcal{B}(\ell)$. Because we need to compute $\pi_{i}$ for each of the $N$ sites, the generation of a new configuration then requires $\mathcal{O}(\mathcal{B}(\ell)\times N)$ steps. We recall that in a $d-$dimensional lattice $\mathcal{B}(\ell)\propto\ell^{d}$, while for random graphs $\mathcal{B}(\ell)\propto c^{\ell}$. It is therefore crucial to assess how the number of layers $\ell$ should scale with the system size $N$.
@@
In this paper, in order to test the 4N architecture on a prototypically hard-to-sample problem, we will consider the Edwards-Anderson spin glass model, described by the Hamiltonian in Eq. (3) for Ising spins, where we set $H_{i}=0$ for simplicity. The couplings are chosen to be non-zero only for neighboring pairs $\langle i,j\rangle$ on a two-dimensional square lattice. Non-zero couplings $J_{ij}$ are independent random variables, identically distributed according to either a normal or a Rademacher distribution. This model was also used as a benchmark in Refs. [13, 12]. All the results reported in the rest of the section are for a $16\times 16$ square lattice, hence with $N=256$ spins, which guarantees a large enough system to investigate correlations at low temperatures while keeping the problem tractable. In the SI we also report some data for a $32\times 32$ lattice.
While this model has no finite-temperature spin glass transition, sampling it via standard local MCMC becomes extremely hard for $\beta\gtrsim 1.5$ [12]. To this day, the best option for sampling this model at lower temperatures is parallel tempering (PT), which we used to construct a sample of $M=2^{16}$ equilibrium configurations for several instances of the model and at several temperatures. Details are given in the Methods section.
@@
We begin by comparing the mean energy of the configurations generated by the model with the mean energy of the configurations generated via parallel tempering, which is important because the energy plays a central role in the Metropolis reweighing scheme of Eq. (5). Fig. 2 shows the relative energy difference, as a function of inverse temperature $\beta$ and number of layers $\ell$, averaged over 10 different instances of the disorder. First, it is clear that adding layers improves results remarkably. We stress that, since the activation functions of our network are linear, adding layers only helps in taking into account spins that are further away. Second, we notice that the relative error found when using 6 layers or more is small for all temperatures considered. Finally, we notice that the behaviour of the error is non-monotonic, and indeed appears to decrease for low temperatures even for networks with few layers.
@@
Figure 4: Examples of the probability distribution of the overlap, $P(q)$, for different instances of the disorder and different temperatures. Solid color lines correspond to configurations generated using 4N, while dashed black lines correspond to the $P(q)$ obtained from equilibrium configurations generated via PT. (a) At high temperature, even a couple of layers are enough to reproduce well the distribution. (b) At lower temperatures more layers are required, as expected because of the increase of the correlation length and of the complexity of the Gibbs-Boltzmann distribution. (c) Even lower temperature for the same instance. (d) A test on a more complex instance shows that 4N is capable to learn even non-trivial forms of $P(q)$.
@@
Figure 6: (a) Comparison between the correlation length $\xi$ and the decay parameter $\overline{\ell}$ of the exponential fit as a function of the inverse temperature $\beta$. Both data sets are averaged over 10 instances. The data are limited to $T\geq 0.45$ because below that temperature the finite size of the sample prevents us to obtain a reliable proper estimate of the correlation length. (b) Absolute difference between $\langle\log P^{\ell}_{\text{NN}}\rangle_{\text{data}}$ at various $\ell$ and at $\ell=10$ for $T=1.02$ and different system sizes. Data are averaged over ten different instances of the disorder and the colored areas identify regions corresponding to plus or minus one standard deviation. A version of this plot in linear-logarithmic scale is available in the SI.
@@
$G(r)=\frac{1}{{\color[rgb]{0,0,0}\mathcal{N}(r)}}\sum_{i,j:d_{ij}=r}\overline{\langle q_{i}q_{j}\rangle}\ ,$ (13)
where $r=d_{ij}$ is the lattice distance of sites $i,j$, ${q_{i}=\sigma^{1}_{i}\sigma^{2}_{i}}$ is the local overlap, and $\mathcal{N}(r)=\sum_{i,j:d_{ij}=r}1$ is the number of pairs of sites at distance $r$. Note that in this way $G(0)=1$ by construction. Results for $G(r)$ at various temperatures are shown in Fig. 5. It is clear that increasing the number of layers leads to a better agreement at lower temperatures. Interestingly, architectures with a lower numbers of layers produce a smaller correlation function at any distance, which shows that upon increasing the number of layers the 4N architecture is effectively learning the correlation of the Gibbs-Boltzmann measure.
To quantify this intuition, we extract from the exact $G(r)$ of the EA model a correlation length $\xi$, from the relation $G(\xi)=1/e$. Then, we compare $\xi$ with $\overline{\ell}$, i.e. the number of layers above which the model becomes accurate, extracted from Fig. 3. These two quantities are compared in Fig. 6 as a function of temperature, showing that they are roughly proportional. Actually $\overline{\ell}$ seems to grow even slower than $\xi$ at low temperatures. In Fig. 6 we also reproduce the results of Fig. 3b for different sizes of the system. The curves coincide for large enough values of $N$, showing that the value of $\ell$ saturates to a system-independent value when increasing the system size. We are thus able to introduce a physical criterion to fix the optimal number of layers needed for the 4N architecture to perform well: $\ell$ must be of the order of the correlation length $\xi$ of the model at that temperature.
We recall that the 4N architecture requires $(c+1)\times\ell\times N$ parameters, where $c$ is the average connectivity of the graph ($c=2d$ for a $d$-dimensional lattice), and $\mathcal{O}(\mathcal{B}(\ell)\times N)$ operations to generate a new configuration. We thus conclude that, for a model with a correlation length $\xi$ that remains finite for $N\to\infty$, we can choose $\ell$ proportional to $\xi$, hence achieving the desired linear scaling with system size both for the number of parameters and the number of operations.